# GUI-World: A Dataset for GUI-oriented Multimodal LLM-based Agents

### v1

@@ Appendix C Details of Experiments Setups
C.1 Finetune dataset construction
We use two settings to finetune GUI-Vid, one with video-text pairs only, and the other with video-text and image-text pairs, which are all GUI content:
Video Only. In this setting, we only trained GUI-Vid with video-text pairs in GUI-World, as shown in Table 9.
Video-Image. Inspired by the pre-trained process of Videochat2, we include image-text pairs to help the visual encoder align GUI knowledge. These images are selected from our GUI-World, MetaGUI [14], and OmniAct [21] for high-quality GUI content. Subsequently, we use GPT-4V to generate a detailed description and a concise caption for each image. Finally, we construct a dataset consisting of video-text and image-text pairs for gaining comprehensive GUI-oriented capabilities.
@@ C.2 Hyperparameter Settings
Inference.
We empirically study 7 MLLMs, involving 4 Image-LLMs and 3 Video-LLMs, with their hyperparameters detailed as follows:
Gemini-Pro-1.5 [42]: We use the default settings, which set temperature as 0.4, top-p as 1, and max-token as 2048. It should be noted that during our project, Gemini-Pro-1.5 is still under the user request limit, which only provides 100 requests per day, making our benchmark difficult. Given that Gemini hasn’t launched Pay-as-you-go444https://ai.google.dev/pricing, we will include benchmark results on ‘Human’ setting as soon as possible.
Qwen-VL-Max [41]: We use the default settings for Qwen-VL-Max, with top-p as 0.8 and max-token as 2048. Given that the input context window is merely 6,000 for Qwen, we scale the resolution for all images to 0.3.
ChatUnivi [44]: We use ChatUnivi-7B built upon Vicuna-v0-7B and set the max frame as 100, temperature as 0.2, and max-token as 1024.
Minigpt4video [45]: We use the suggested settings555https://github.com/Vision-CAIR/MiniGPT4-video for this model and the max-frame are set as 45, with only the max-token being modified to 1024.
VideoChat2 & GUI-Vid [46]: For a fair comparison, we set the same hyperparameters for VideoChat2 & GUI-Vid. We set the max-token as 1024, top-p as 0.9, temperature as 1.0, max-frame as 8/16, repetition penalty as 1.2, and length penalty as 1.2.
LLM-as-a-Judge.
We studied four LLM-as-a-Judge in giving a similarity score for the MLLM’s response and ground truth, namely GPT-4 [52], ChatGPT [96], LLaMA-3-70b-instruct [54], and Mixtral-8x22b-instruct-v0.1 [55]. Hyperparameter settings are detailed as follows:
@@
Finetune.
We include several hyperparameter settings in experiment settings and ablation studies, as shown in Table 11.
@@
Given the complexity of free-form answers in GUI scenarios, the evaluation includes specific positions of GUI elements, textual content, and comparing the response to the golden answer. LLM-as-a-judge has been widely used in previous studies for complex evaluation tasks [47, 48]. Therefore, we leverage LLM-as-a-Judge [47] in a similar setting to MM-vet [60], which compares the MLLM’s response to the golden answer. We carefully evaluate the accessibility of leveraging LLM-as-a-Judge, selecting 1,000 samples covering 6 free-form questions mentioned in our dataset. As shown in Table 12, GPT-4 outperforms other LLMs, exhibiting a better human alignment on providing a similarity score for the response compared to the golden answer, although it is approximately 10 times more expensive than other models.
@@ Appendix D Additional Experiments Results
In this section, we provide detailed results on each task in each GUI scenario. For captioning tasks, Table 13 shows comprehensive experimental results among six scenarios. For scores of LLM-as-a-Judge in a specific task, see Table 14, Table 15, Table 16, Table 17, and Table 18. For BLEU [50] and BERTScore [51] in validating free-form and conversational questions, see Table 19, Table 20, Table 21, Table 24, Table 22, and Table 23. For performance in fine-grain (application level), see Figure 16 for Gemini-Pro and Figure 17 for Qwen-VL-Max.

### v2

@@ Appendix C Details of Experiments Setups
C.1 Fine-tune dataset construction
We use two settings to fine-tune GUI-Vid, one with video-text pairs only, and the other with video-text and image-text pairs, which are all GUI content:
Video Only. In this setting, we only train GUI-Vid with video-text pairs in GUI-World, as shown in Table 12.
Video-Image. Inspired by the pre-trained process of Videochat2, we include image-text pairs to help the visual encoder align GUI knowledge. These images are selected from our GUI-World, MetaGUI (Sun et al., 2022), and OmniAct (Kapoor et al., 2024) for high-quality GUI content. Subsequently, we use GPT-4V to generate a detailed description and a concise caption for each image. Finally, we construct a dataset consisting of video-text and image-text pairs for gaining comprehensive GUI-oriented capabilities.
@@ C.2 Hyperparameter Settings
Inference. We empirically study 7 MLLMs, involving 4 Image LLMs and 3 Video LLMs, with their hyperparameters detailed as follows:
Gemini-Pro-1.5 (Team et al., 2023): We use the default settings, which set temperature as 0.4, top-p as 1, and max-token as 2048. It should be noted that during our project, Gemini-Pro-1.5 is still under the user request limit, which only provides 100 requests per day, making our benchmark difficult. Given that Gemini hasn’t launched Pay-as-you-go444https://ai.google.dev/pricing, we will include benchmark results on ‘Human’ setting as soon as possible.
Qwen-VL-Max (Bai et al., 2023): We use the default settings for Qwen-VL-Max, with top-p as 0.8 and max-token as 2048. Given that the input context window is merely 6,000 for Qwen, we scale the resolution for all images to 0.3.
ChatUnivi (Jin et al., 2023): We use ChatUnivi-7B built upon Vicuna-v0-7B and set the max frame as 100, temperature as 0.2, and max-token as 1024.
Minigpt4video (Ataallah et al., 2024): We use the suggested settings555https://github.com/Vision-CAIR/MiniGPT4-video for this model and the max-frame are set as 45, with only the max-token being modified to 1024.
VideoChat2 & GUI-Vid (Li et al., 2023c): For a fair comparison, we set the same hyperparameters for VideoChat2 & GUI-Vid. We set the max-token as 1024, top- as 0.9, temperature as 1.0, max-frame as 8/16, repetition penalty as 1.2, and length penalty as 1.2.
LLM-as-a-Judge. We investigate four LLM-as-a-Judge in giving a similarity score for the MLLM’s response and ground truth, namely GPT-4 (Achiam et al., 2023), ChatGPT (OpenAI, 2023), LLaMA-3-70b-instruct (Meta, 2023b), and Mixtral-8x22b-instruct-v0.1 (ai, 2024). Hyperparameter settings are detailed as follows:
@@
Fine-tune. We include several hyperparameter settings in experiment settings and ablation studies, as shown in Table 14.
@@
Given the complexity of free-form answers in GUI scenarios, the evaluation includes specific positions of GUI elements, textual content, and comparing the response to the golden answer. LLM-as-a-Judge has been widely used in previous studies for complex evaluation tasks (Zheng et al., 2023, Liu et al., 2023d). Therefore, we leverage LLM-as-a-Judge (Zheng et al., 2023) in a similar setting to MM-vet (Yu et al., ), which compares the MLLM’s response to the golden answer. We carefully evaluate the accessibility of leveraging LLM-as-a-Judge, selecting 1,000 samples covering 6 free-form questions mentioned in our dataset. As shown in Table 15, GPT-4 outperforms other LLMs, exhibiting a better human alignment on providing a similarity score for the response compared to the golden answer, although it is approximately 10 times more expensive than other models.
@@ Appendix D Additional Experiments Results
Removing Data from MetaGUI and OmniAct. As shown in Table 16, after excluding OmniAct and MetaGUI from the training data, the model showed performance decreases in certain areas while demonstrating notable improvements in others. Specifically, we observed significant enhancements in MCQA tasks and XR-related capabilities. Conversely, models trained with MetaGUI and OmniAct exhibited superior performance on general free-form questions. These experimental results validate that models trained exclusively on GUI-World can achieve highly competitive performance, thereby confirming the exceptional quality of our dataset.
| Models | Software | Website | XR | Multi | IOS | Android | Avg. | |||||||||
| MC | Free | MC | Free | MC | Free | MC | Free | MC | Free | MC | Free | MC | Free | |||
| VideoChat2 | 45.50% | 2.144 | 42.60% | 2.221 | 44.00% | 2.005 | 40.40% | 2.222 | 40.20% | 2.169 | 44.70% | 2.119 | 42.90% | 2.147 | ||
| GUI-Vid | 59.90% | 2.847 | 54.10% | 2.957 | 55.60% | 2.764 | 52.90% | 2.861 | 51.80% | 2.773 | 53.40% | 2.572 | 54.60% | 2.796 | ||
59.70% | 2.811 | 57.30% | 2.844 | 59.62% | 2.853 | 50.00% | 2.799 | 55.31% | 2.769 | 57.67% | 2.547 | 56.60% | 2.767 | ||
| Scenarios | GUI-Vid | Tie | VideoChat2 |
| Software | 82.7% | 13.3% | 4.0% |
| Website | 86.0% | 12.0% | 2.0% |
| XR | 88.0% | 8.7% | 3.3% |
| Multi | 85.3% | 10.0% | 8.7% |
| IOS | 92.0% | 6.0% | 2.0% |
| Android | 82.0% | 16.0% | 2.0% |
| Average | 86.0% | 11.0% | 3.7% |
| Model | GUI-World | VisualAgentBench | VideoGUI | OS-World |
| GPT-4o | 1 | 1 | 1 | 2 |
| GPT-4V | 2 | 2 | 2 | 1 |
| Gemini-1.5-Pro | 3 | 3 | 3 | 3 |
| Qwen-VL-Max | 4 | 4 | 4 | / |
In this section, we first provide an ablation study on keyframe selection methods. Then, we conduct statistics and human preference experiments on correlations of GUI understanding capability to other mainstream GUI-related tasks. Furthermore, we provide detailed, performance on newly released models after our submission of the first version, followed by very detailed results on each task in each GUI scenario.
Ablation Study on Keyframe Identify Methods. Firstly, we show performance on model-based keyframe identify methods in Table 6, with details of UVD+VIP and UVD+R3M in Tables 24 and 25.
Correlation between GUI Understanding and Other Mainstream GUI Tasks. Furthermore, we conduct additional analysis and experiments to show how GUI understanding capability helps mainstream GUI-related tasks, including generating code to operate GUI (Cheng et al., 2024) and assist people through chat (Hong et al., 2024). Both demonstrate the strong correlation between GUI understanding capability and specific tasks for GUI agents.
We compare the benchmark results on GUI-World with existing benchmarks (Xie et al., 2024, Qinghong Lin et al., 2024, Liu et al., 2024c) for operating on GUI as shown in Table 18, and find that the results generally match, i.e., the stronger the understanding ability, the stronger the agent performance.
For the definition of chat helping humans, we select 180 videos from the benchmark, choosing 30 videos for each scenario. We ask 5 human annotators to pose the question they most wanted to ask after watching each video. We then use GUI-Vid, both before and after fine-tuning, to answer these questions. The human annotators who ask the questions are then asked to indicate which answer is more helpful. The results are shown in Table 17, demonstrating that models trained in GUI understanding are more favored by people when acting as GUI agents.
| Scenario | MCQA | Description | Conversation | Dynamic | Static | Caption | Average |
| XR | 0.442 | 1.100 | 2.686 | 2.055 | 1.808 | 1.654 | 2.258 |
| Android | 0.513 | 1.162 | 2.952 | 1.858 | 1.673 | 1.763 | 2.259 |
| IOS | 0.497 | 1.143 | 2.966 | 1.992 | 1.680 | 1.654 | 2.319 |
| Multi | 0.459 | 1.106 | 2.863 | 2.069 | 1.781 | 1.772 | 2.329 |
| Website | 0.524 | 1.183 | 3.059 | 2.102 | 1.736 | 1.371 | 2.410 |
| Software | 0.529 | 1.241 | 2.942 | 1.958 | 1.657 | 1.519 | 2.290 |
| Average | 0.494 | 1.156 | 2.911 | 2.005 | 1.722 | 1.622 | 2.311 |
| Scenarios | MCQA | Description | Conversation | Dynamic | Static | Caption | Average |
| XR | 0.596 | 1.867 | 3.123 | 2.580 | 2.147 | 1.987 | 2.709 |
| Android | 0.243 | 1.675 | 3.338 | 2.360 | 1.980 | 2.189 | 2.675 |
| IOS | 0.581 | 1.762 | 3.229 | 2.536 | 2.051 | 2.017 | 2.717 |
| Multi | 0.355 | 1.069 | 2.982 | 2.437 | 1.870 | 2.541 | 2.541 |
| Website | 0.484 | 1.729 | 3.123 | 2.422 | 1.854 | 2.004 | 2.588 |
| Software | 0.569 | 1.762 | 3.220 | 2.448 | 1.868 | 2.149 | 2.641 |
| Average | 0.471 | 1.644 | 3.169 | 2.464 | 1.961 | 2.148 | 2.645 |
Performance of Newly Released Models in GUI-World Test Set. We evaluate two latest models, LLaVA-Next-Video-7B-DPO (Liu et al., 2024a) and Video-LLaVA (Lin et al., 2023a). We show their performance in Tables 20 and 19. Our model outperforms these in most tasks, except conversation, likely due to their use of DPO during training.
Performance on MV-Bench. We conducted evaluations on MVBench (Li et al., 2023c) across four models, as shown in Table 21. While our model showed some performance degradation compared to the original model, it still significantly outperformed the other two baseline models. We would like to note that the primary purpose of training GUI-Vid was to demonstrate the effectiveness of our dataset, rather than achieving state-of-the-art performance. Therefore, we did not employ the strongest baseline model or optimal training methodologies, which may have contributed to some catastrophic forgetting. However, given that GUI-Vid is specifically designed for GUI-related tasks, we believe this performance trade-off is acceptable within our research context.
| Category | VideoChat2 | GUI-Vid | VideoLLaMA | VideoChatGPT |
| Action Prediction | 47.5 | 39.0 | 25.5 | 26.0 |
| Unexpected Action | 60.0 | 57.0 | 39.0 | 26.5 |
| Object Existence | 58.0 | 54.0 | 48.0 | 54.0 |
| Object Interaction | 71.5 | 51.0 | 40.5 | 28.0 |
| Object Shuffle | 41.0 | 30.5 | 38.0 | 40.0 |
| Moving Direction | 23.0 | 21.0 | 22.5 | 23.0 |
| Action Localization | 23.0 | 30.5 | 22.5 | 20.0 |
| Scene Transition | 88.0 | 65.5 | 43.0 | 31.0 |
| Action Count | 39.5 | 34.5 | 34.0 | 30.5 |
| Moving Count | 42.0 | 29.0 | 22.5 | 25.5 |
| Moving Attribute | 58.5 | 53.5 | 32.5 | 39.5 |
| State Change | 44.5 | 41.0 | 45.5 | 48.5 |
| Character Order | 36.5 | 39.0 | 32.5 | 29.0 |
| Egocentric Navigation | 35.0 | 34.0 | 40.0 | 33.0 |
| Episodic Reasoning | 38.5 | 37.0 | 30.0 | 29.5 |
| Average | 47.1 | 41.1 | 34.4 | 32.3 |
Performance on Mind2Web-Multimodal. We evaluated our model on the Mind2Web-Multimodal dataset (Deng et al., 2024) to demonstrate the effectiveness of fine-tuning on GUI-World. Mind2Web-Multimodal is a multiple-choice GUI benchmark that assesses GUI operation capabilities, where each sample comprises an image, a task description, and response options. Our evaluation included four experimental settings: zero-shot inference using both VideoChat2 and GUI-Vid, as well as versions of these models fine-tuned on Mind2Web-Multimodal. We conducted the fine-tuning process for 3 epochs. As shown in Tables 22 and 23, fine-tuning on GUI-World enhances the models’ GUI understanding capabilities and improves their performance on operational benchmarks.
| Category | ZeroShot | Fine-tuned | ||
| VideoChat2 | GUI-Vid | VideoChat2 | GUI-Vid | |
| Cooking | 21.28% | 23.40% | 23.40% | 26.24% |
| Education | 17.61% | 16.61% | 24.92% | 28.24% |
| Finance | 26.02% | 29.59% | 26.02% | 32.65% |
| Government | 18.80% | 18.28% | 26.89% | 27.42% |
| Health | 17.08% | 17.37% | 22.58% | 26.48% |
| Home service | 19.08% | 20.14% | 21.55% | 28.27% |
| Housing | 12.21% | 14.13% | 18.20% | 21.84% |
| Job | 16.25% | 18.00% | 20.00% | 20.25% |
| Moving | 14.36% | 17.44% | 20.00% | 24.62% |
| Pet | 14.05% | 17.97% | 21.57% | 21.24% |
| Shipping | 13.42% | 12.99% | 20.78% | 20.34% |
| Social media | 22.15% | 24.37% | 22.78% | 29.43% |
| Weather | 27.14% | 22.14% | 26.43% | 20.00% |
| Overall | 17.53% | 18.59% | 22.37% | 25.14% |
| Category | ZeroShot | Fine-tuned | ||
| VideoChat2 | GUI-Vid | VideoChat2 | GUI-Vid | |
| Auto | 20.00% | 16.00% | 19.00% | 18.00% |
| Department | 9.52% | 11.90% | 14.29% | 16.67% |
| Digital | 19.73% | 19.73% | 15.65% | 20.41% |
| Event | 34.65% | 25.74% | 21.78% | 22.77% |
| General | 18.32% | 18.85% | 17.80% | 20.94% |
| Music | 16.67% | 15.66% | 21.69% | 24.10% |
| Other | 17.81% | 21.92% | 27.40% | 35.62% |
| Restaurant | 10.70% | 14.44% | 17.65% | 16.58% |
| Sports | 16.98% | 18.87% | 20.75% | 26.42% |
| Overall | 17.94% | 17.96% | 18.84% | 21.20% |
@@
Additional Experiment Results on GUI-World. For captioning tasks, Table 26 shows comprehensive experimental results among six scenarios. For scores of LLM-as-a-Judge in a specific task, see Tables 27, 28, 29, 30, and 31. For performance in fine-grain (application level), see Figure 16 for Gemini-Pro and Figure 17 for Qwen-VL-Max.
| Scenario | MCQA | Description | Conversation | Dynamic | Static | Caption | Average |
| Software | 86.2% | 3.297 | 4.282 | 3.354 | 3.478 | 4.112 | 3.749 |
| Website | 82.0% | 3.248 | 4.155 | 3.415 | 3.567 | 4.074 | 3.744 |
| XR | 84.2% | 2.980 | 3.775 | 3.034 | 3.122 | 3.587 | 3.347 |
| Multi | 82.1% | 3.391 | 4.165 | 3.466 | 3.404 | 3.868 | 3.659 |
| IOS | 86.0% | 3.157 | 4.017 | 3.353 | 3.492 | 4.050 | 3.648 |
| Mobile | 80.7% | 2.827 | 3.871 | 2.970 | 3.014 | 3.844 | 3.340 |
| Average | 83.5% | 3.150 | 4.044 | 3.265 | 3.346 | 3.923 | 3.581 |
| Scenario | MCQA | Description | Conversation | Dynamic | Static | Caption | Average |
| Software | 85.8% | 3.290 | 4.273 | 3.352 | 3.458 | 4.134 | 3.741 |
| Website | 82.7% | 3.282 | 4.114 | 3.460 | 3.591 | 4.065 | 3.746 |
| XR | 87.7% | 3.010 | 3.861 | 3.142 | 3.161 | 3.600 | 3.433 |
| Multi | 83.6% | 3.237 | 4.129 | 3.503 | 3.417 | 3.897 | 3.737 |
| IOS | 86.4% | 3.165 | 4.094 | 3.328 | 3.480 | 4.078 | 3.663 |
| Android | 80.6% | 2.835 | 3.876 | 2.968 | 3.072 | 3.865 | 3.353 |
| Average | 84.5% | 3.136 | 4.058 | 3.292 | 3.363 | 3.940 | 3.612 |